OpenStudy (anonymous):

Verify the identity cot^2x + csc^2x = 2csc^2x - 1

OpenStudy (anonymous):

please help me

OpenStudy (mertsj):

\[\frac{\cos ^2x}{\sin ^2x}+\frac{1}{\sin ^2x}\] \[\frac{1-\sin ^2x}{\sin ^2x}+\frac{1}{\sin ^2x}\] \[\frac{2-\sin ^2x}{\sin ^2x}\] \[\frac{2}{\sin ^2x}-\frac{\sin ^2x}{\sin ^2x}\] \[2\csc ^2x-1\]

OpenStudy (mertsj):

Don't you just hate it when someone posts a question and then goes offline to wait for someone else to do their work?

OpenStudy (anonymous):

nono no im so sorry @Mertsj i had to call my father see im not living with him any more and him bein 3000 miles away he likes to call alot i wasnt trying to go offline to wait i wanted to work with you on it for a thourough explanation i dont just want the work i want to know how its done

OpenStudy (anonymous):

@Mertsj